OpenStudy (younginslim):

What is the equation, in slope-intercept form, of the line that passes through (0, 5) and has a slope of −1? y = −x − 5 y = x + 5 y = −x + 5 y = x − 5

OpenStudy (visarika23):

Equation of line in slope-intercept form is (y-y1)=(x-x1)*m where x1 and y1 are the points through which the line passes and m is the slope

OpenStudy (visarika23):

In this q y1=5 and x1=0 and m=-1 (y-5)=(x-0)*(-1) y=-x+5